Morning Litho status report 2/25

Gary Sosa gsosa at stanford.edu
Fri Feb 25 07:31:12 PST 2011


Good morning All.....

    Here is the morning Litho status report.

EV Aligner- The lamp house optics are being replaced and 
intensity/uniformity tested

EV Bonder- Repairs and testing of heater lamps to be completed 
today(2/25). The bonder will be configured for "Steel"
                       pressure plate as per the published schedule 
posted at the machine.

Dump Rinser- User reports problem with D/R in "wbmiscres". Will test out 
this morning

Microscope 2- Across from Drytek. User reports a problem. Will check out 
microscope this morning.

     Please check back in Coral for latest tool updates and current 
equipment status.

       Thanks...  The Litho Team


*/Photolithography Area Status/*

	

	

*/Friday, February 25, 2011/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

Mahnaz- Out today 2/25

*/Exposure Tools/**//*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*UP*

	

	

EVAlign

	

*DOWN*

	

Replacing optics

	

2

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*UP*

	

	

Karl Suss 2

	

*UP*

	

	

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*UP*

	

	

Headway

	

*UP*

	

	

Laurel

	

*UP*

	

	

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Up*

	

	

220-3 Process

	

*Up*

	

	

220-7 Process

	

*Problem*

	

Optimizing Process

	

4

SVGCoat2

	

3612 Process

	

*Up*

	

	

955-0.7 Process

	

*Up*

	

	

3617 Process

	

*Up*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*UP*

	

	

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*UP*

	

	

Blue M 150 C

	

*UP*

	

	

Fusion

	

*DOWN*

	

Hotplate experiments

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*UP*

	

	

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*PROBLEM*

	

-11 stage error

	

5

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*PROBLEM*

	

Not working correctly

	

3

#3 Olympus/Litho

	

*UP*

	

	

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*DOWN*

	

Top heater not working

	

1

KSBonder

	

*PROBLEM*

	

Leaking nitrogen hose

	

6

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

Mask Cleaner

	

*UP*

	

	

	

	

	


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20110225/0a9a1df1/attachment.html>


More information about the litho mailing list